### v1

@@ 4.1 Settings
Grok-4 \$3.98 \$11.11 \$8.97 \$171.56 \$45.78 \$2.69 \$50.08 \$22.44 \$29.24 \$345.85 ($\downarrow$ 81.68%) 64.17
GPT-5 \$1.57 \$7.97 \$2.22 \$39.29 \$18.17 \$1.43 \$10.25 \$16.98 \$7.49 \$105.37 ($\downarrow$ 39.87%) 64.30
GPT-5.2-Thinking \$1.07 \$5.39 \$1.64 \$52.49 \$15.62 \$1.18 \$6.49 \$16.80 \$11.52 \$112.20 ($\downarrow$ 43.53%) 66.20
Gemini-3-Pro \$1.73 \$4.57 \$3.18 \$52.47 \$22.91 \$1.97 \$23.29 \$7.56 \$17.78 \$135.46 ($\downarrow$ 53.23%) 71.00
@@
4.2 Main Results
As demonstrated in Table 1, the proposed JiSi is compared with the aforementioned closed-source and open-source LLMs, along with typical router methods and multi-agent methods, i.e., RouterDC (chen2024routerdc), EmbedLLM (zhuang2024embedllmlearningcompactrepresentations), GraphRouter (feng2025graphroutergraphbasedrouterllm), Avengers (zhang2025avengerssimplerecipeuniting), Retrieval-based Router (tang2025opensourcellmscollaborationbeats), MOA (wang2024mixtureofagentsenhanceslargelanguage), Self Consistency (chen2023universalselfconsistencylargelanguage), and Majority Voting (chen2024llmcallsneedscaling). For multi-agent baselines, priors from OpenRouterBench are applied to provide a stronger variant. For each benchmark, the MOA selects three LLMs with the highest evaluation scores over the training records on the benchmark; the Self Consistency sample four rollouts from the best LLM evaluated on training data and merges responses via majority voting; the Majority Voting is verified in two settings. The first is vanilla, encompassing all open-source LLMs, whereas the second comprises only the top-3 LLMs based on the training set. Besides, the best result of open-source LLMs on the test set of each benchmark is gathered to form the model entry “Best LLM” in Table 1.
Superiority of JiSi. The proposed JiSi framework exhibits promising performance, ranking first on average among all closed-source LLMs, open-source LLMs, router methods, and multi-agent methods. The performance gain over the most powerful open-source LLM (DeepSeek-V3.2-Speciale) is substantially 6.56%. Even compared with the state-of-the-art closed-source LLM Gemini-3-Pro, a +1.15% gain is achieved. Moreover, the router-only variant of JiSi, which directly adopts $\boldsymbol{A}_{1}^{fm}$ without the adaptive aggregation formulated in Eq. 5, outperforms other router methods, indicating that the proposed query-response mixed router can yield more suitable aggregatees for subsequent aggregation. Compared with the router-only variant, JiSi (w/o Adaptive Aggregation) benefits +1.41% from dynamic aggregator selection, and the proposed adaptive aggregation further improves the performance by +1.06%. Besides, with the additional aggregation, JiSi overpasses the theoretical performance bound of routing, i.e. “Best LLM”, by +1.6%, revealing enormous potential.
Efficiency of JiSi. As shown in Table 2, we report the cost of the proposed JiSi and the leading proprietary LLMs. It is worth noting that although JiSi utilizes more API calls due to multi-agent aggregation, it can remarkably reduce the total cost while maintaining superior performance compared with leading proprietary LLMs. For instance, JiSi surpasses Gemini-3-Pro and Grok-4 by +1.15% and +7.89%, while saving 53.23% and 81.68% cost. The reasons lie in two aspects: 1) the APIs of open-source LLMs are dramatically cheaper than proprietary LLMs. 2) The proposed adaptive routing-aggregation switch can mitigate the workload of aggregation and even simplify aggregation to routing. The strong trade-off between efficiency and effectiveness demonstrates the feasibility and economy of JiSi in practical implementation.
@@ 4.2 Main Results
Is Self-Consistency Good Enough? The answer is “No”. It is worth noting that the multi-agent baseline “Self Consistency” (SC) seems to be performant enough, almost reaching the performance of the best closed-source LLM with a gap within 0.5%. However, the strong results from SC are based on an “oracle” setting that assumes the method can determine to which benchmark a new test question belongs and apply the most scored LLMs evaluated on the training set of this benchmark. It should be clarified that such a setting is practically infeasible, as the type of a given question is generally unknown. In contrast, the proposed JiSi is question-type-agnostic, using similarity metrics to construct a support set for LLM routing, which is practical and extensible. Moreover, the four-rollout SC, which costs almost equally as JiSi, does not provide a significant boost (just about 0.1%) compared with “Best LLM”, demonstrating the limitation of homogeneous LLM scaling.
@@
Compared with Voting. Majority voting is widely used to make multi-agent decisions. However, it is fragile to LLM responses, especially when most collaborative LLMs cannot correctly handle a question. Voting may be biased toward a majority of incorrect answers, leading to worse results when scaling, as shown in Table 1 (-2.95% from 3 “Best LLMs” to “10 LLMs”). The proposed JiSi explores the minority of correct answers via fine-grained rerouting, thereby ensuring scalability while delivering extraordinary performance.
@@ 4.3 Scaling Ability
To verify the scalability of JiSi, we initialize a base LLM bank with five LLMs: DeepSeek-V3-0324, DeepSeek-R1-0528, Intern-S1, Qwen3-235B-A22B-2507, and Qwen3-235B-A22B-Thinking-2507. Subsequently, we introduce five additional open-source LLMs following their chronological release sequence: Kimi-K2-0905, DeepSeek-V3.1-Terminus, GLM-4.6, DeepSeek-V3.2-Thinking, and DeepSeek-V3.2-Speciale. This setup allows us to simulate the real-world evolution of the LLM ecosystem, resulting in five experimental configurations with model pool sizes $\{5,6,7,8,9,10\}$. All other hyperparameters remain the same as Sec. 4.1. To provide a multi-faceted evaluation, we categorize nine benchmarks into three specialized domains: 1) Reasoning & Knowledge: AIME, GPQA, MMLU-Pro, HLE, LiveMathBench. 2) Coding & Engineering: LiveCodeBench, SWE-bench. 3) General Chat & Factuality: Arena-Hard, SimpleQA. As illustrated in Fig. 5, JiSi demonstrates a robust and consistent performance gain across all three domains as the number of candidate LLMs increases. The consistent improvement across diverse benchmarks validates that JiSi is not merely a static ensemble but a dynamic system that scales with the progress of the open-source community. This empirical evidence supports a paradigm shift in intelligence enhancement, transitioning from the neural scaling law of parameters and training data towards a novel scaling law of multi-agent collaboration.
4.4 Distribution of LLM Selection

### v2

@@ 4.2 Main Results
Grok-4 (xAI, 2025) \$3.98 \$11.11 \$8.97 \$171.56 \$45.78 \$2.69 \$50.08 \$22.44 \$29.24 \$345.85 ($\downarrow$ 81.68%) 64.17
GPT-5 (OpenAI, 2025b) \$1.57 \$7.97 \$2.22 \$39.29 \$18.17 \$1.43 \$10.25 \$16.98 \$7.49 \$105.37 ($\downarrow$ 39.87%) 64.30
GPT-5.2-Thinking (OpenAI, 2025a) \$1.07 \$5.39 \$1.64 \$52.49 \$15.62 \$1.18 \$6.49 \$16.80 \$11.52 \$112.20 ($\downarrow$ 43.53%) 66.20
Gemini-3-Pro (Google, 2025) \$1.73 \$4.57 \$3.18 \$52.47 \$22.91 \$1.97 \$23.29 \$7.56 \$17.78 \$135.46 ($\downarrow$ 53.23%) 71.00
@@
Is Self-Consistency Good Enough? The answer is “No”. It is worth noting that the multi-LLM baseline “Self Consistency” (SC) seems to be performant enough, almost reaching the performance of the best closed-source LLM with a gap within 0.5%. However, the strong results from SC are based on an “oracle” setting that assumes the method can determine to which benchmark a new test question belongs and apply the most scored LLMs evaluated on the training set of this benchmark. It should be clarified that such a setting is practically infeasible, as the type of a given question is generally unknown. In contrast, the proposed JiSi is question-type-agnostic, using similarity metrics to construct a support set for LLM routing, which is practical and extensible. Moreover, the four-rollout SC, which costs almost equally as JiSi, does not provide a significant boost (just about 0.1%) compared with “Best LLM”, demonstrating the limitation of homogeneous LLM scaling.
@@
Compared with Voting. Majority voting (Chen et al., 2025a, 2024a) is widely used to make multi-LLM decisions. However, it is fragile to LLM responses, especially when most collaborative LLMs cannot correctly handle a question. Voting may be biased toward a majority of incorrect answers, leading to worse results when scaling, as shown in Table 1 (-2.95% from 3 “Best LLMs” to “10 LLMs”). The proposed JiSi explores the minority of correct answers via fine-grained rerouting, thereby ensuring scalability while delivering extraordinary performance.
Table 3: The deviation of model performance estimation via different similarity criteria. $\downarrow$: The less is the better.
AIME $\downarrow$ Arena-Hard $\downarrow$ GPQA $\downarrow$ HLE $\downarrow$ LiveCodeBench $\downarrow$ LiveMathBench $\downarrow$ MMLU-Pro $\downarrow$ SimpleQA $\downarrow$ SWE-bench $\downarrow$ Avg $\downarrow$
Query 36.42 35.96 35.18 22.33 37.93 39.72 30.01 37.76 34.31 34.40
Query+Response 35.26 34.10 33.43 22.04 33.29 38.85 27.84 37.00 34.80 32.96
Query+Response+Length 30.92 33.91 31.11 22.51 28.30 37.60 25.40 37.11 34.53 31.27
@@ 4.3 Scaling Ability
To verify the scalability of JiSi, we initialize a base LLM bank with five LLMs: DeepSeek-V3-0324, DeepSeek-R1-0528, Intern-S1, Qwen3-235B-A22B-2507, and Qwen3-235B-A22B-Thinking-2507. Subsequently, we introduce five additional open-source LLMs following their chronological release sequence: Kimi-K2-0905, DeepSeek-V3.1-Terminus, GLM-4.6, DeepSeek-V3.2-Thinking, and DeepSeek-V3.2-Speciale. This setup allows us to simulate the real-world evolution of the LLM ecosystem, resulting in five experimental configurations with model pool sizes $\{5,6,7,8,9,10\}$. All other hyperparameters remain the same as Sec. 4.1. To provide a multi-faceted evaluation, we categorize nine benchmarks into three specialized domains: 1) Reasoning & Knowledge: AIME, GPQA, MMLU-Pro, HLE, LiveMathBench. 2) Coding & Engineering: LiveCodeBench, SWE-bench. 3) General Chat & Factuality: Arena-Hard, SimpleQA. As illustrated in Fig. 5, JiSi demonstrates a robust and consistent performance gain across all three domains as the number of candidate LLMs increases. The consistent improvement across diverse benchmarks validates that JiSi is not merely a static ensemble but a dynamic system that scales with the progress of the open-source community. This empirical evidence supports a paradigm shift in intelligence enhancement, transitioning from the neural scaling law of parameters and training data towards a novel scaling law of multi-LLM collaboration. Moreover, we also demonstrate the task scaling ability of JiSi in Apendix B.
4.4 Analysis on Similarity Evaluation
To evaluate the effectiveness of the proposed query-response mixed similarity discussed in Sec. 3, we propose a new metric. For a given query $x$, the support set can be constructed using various similarity criteria: query similarity only, query with response similarity, and the method adopted by JiSi, which incorporates query, response, and length (token costs) similarity. Once the support set is obtained, a fine-grained model score $\boldsymbol{g}$ can be computed in accordance with Eq. 3. This predicted model score can then be compared to the actual label $\boldsymbol{v}$ using the formula $100\left|\boldsymbol{g}-\boldsymbol{v}\right|_{1}/M$. This represents the deviation between the optimal support set and the actual selected set, and the less the better. In Table 3, this metric is summarized for each benchmark based on the three similarity criteria mentioned earlier. It is evident that, by adding response and length similarity, the deviations are reduced by 1.44% and 1.69%, respectively. This indicates that the proposed mixed-similarity approach provides a more accurate estimate of model performance, causing more reliable routing results for subsequent aggregation.
4.5 Reference Accuracy Study
The proposed Adaptive Routing-Aggregation Switch is designed to filter out ineffective aggregatees while retaining valuable ones. In Table 4, we present the effects of this switching mechanism. To perform a comprehensive evaluation, the embedding bank is divided into three categories based on question difficulty. The detailed division metric is shown in Appendix D.5. The results indicate that the switching mechanism consistently improves aggregatee quality by eliminating less effective ones, particularly for “Hard” questions, where we observe an accuracy increase of 8.49%. This improvement in aggregator performance ensures reliable aggregation by yielding more correct responses.
4.6 Aggregator Selection Study
To validate the aggregator selection rethinking in Sec. 3, we fix three reference responses for each test query while varying the size of the support set to select the aggregator. As the support set size increases, the aggregator exhibits diminished domain-specific capability but enhanced comprehensive capability. As shown in Fig. 6, the aggregator yields suboptimal performance when the aggregator leans excessively towards either domain-specificity or comprehensiveness. The optimal performance is achieved when these two capabilities are well-balanced. It reveals the mechanism behind the proposed Support-Set-based Aggregator Selection, which aims to strike a balance between two capabilities.
Figure 6: The performance curve of aggregators with different domain-specific and comprehensive abilities.
Table 4: The accuracy(%) of 1) all 2) removed 3) remained reference responses when applying adaptive switch.
All Ref Acc Removed Ref Acc Remained Ref Acc
Easy 95.37 93.95(-1.42) 98.07(+2.7)
Medium 71.97 67.54(-4.43) 77.11(+5.14)
Hard 22.25 16.46(-5.79) 30.74(+8.49)
5 Conclusion
We propose JiSi, a training-free and scalable multi-LLM system that leverages a minimalist “Route-and-Aggregate” topology. By incorporating three innovations, including query-response mixed routing, support-set-based aggregator selection, and adaptive routing-aggregation switch, JiSi orchestrates the collaboration of open-source LLMs while mitigating noise. Experiments across nine benchmarks demonstrate that JiSi not only significantly outperforms existing multi-LLM baselines but also surpasses the leading proprietary model, Gemini-3-Pro, while saving 53.23% cost. It verifies that the strategic collaboration of models provides an effective and scalable alternative to the continuous expansion of monolithic models.
Appendix A Distribution of LLM Selection